OpenStudy (anonymous):

Water freezes at 0 degrees celcius and 32 degrees fahrenheit water boils at 100 degrees celcius and 212 degrees fahrenheit, write celsius as a linear function of degrees fahrenheit

OpenStudy (amistre64):

well, they give you 2 points for reference :)

OpenStudy (amistre64):

32,0 and 212,100 determine the slope between them and calibrate it

OpenStudy (anonymous):

Let C = aF+b So 0=32a+b and 100 = 212a+b Solve and simplify for a and b

OpenStudy (amistre64):

100/180 is the slope; 5/9 5/9F -(5(9)/9) + 100 C = (5/9)F +100 maybe?

OpenStudy (anonymous):

should i do y1-y2/x1-x2=slope

OpenStudy (amistre64):

that would work too

OpenStudy (amistre64):

100 - 0 ------- = slope yep 212 - 32